# ReactDance: Hierarchical Representation for High-Fidelity and Coherent Long-Form Reactive Dance Generation

### v2

@@ 4.1 DATASETS AND EXPERIMENTAL SETTING
Evaluation Metrics. We adopt a comprehensive set of quantitative metrics to evaluate the generated motion from four key perspectives: (1) Motion Quality. To assess the realism of individual motions, we measure distributional similarity and reconstruction accuracy. Distributional metrics include Fréchet Inception Distance (FID) and Diversity (Div) on both kinematic ($\text{FID}_{k},\text{Div}_{k}$(Onuma et al., 2008) and graphical ($\text{FID}_{g},\text{Div}_{g}$(Müller et al., 2005) features. Reconstruction accuracy against the ground truth is measured by Mean Per-Joint Position Error (MPJPE) for static poses and Mean Per-Joint Velocity Error (MPJVE) for dynamics (Ghosh et al., 2024). (2) Interaction Quality. To evaluate leader-reactor coordination, we compute FID and Diversity on cross-distance features ($\text{FID}_{cd},\text{Div}_{cd}$), which describe the evolving spatial relationships between the dancers’ key joints. (3) Music-Dance Alignment. We use two metrics: the Beat Align Score (BAS) (Siyao et al., 2022) to measure synchronization between motion peaks and music beats, and the Beat Echo Degree (BED) (Siyao et al., 2024) to quantify rhythmic consistency between the dancers. (4) Generation Efficiency. We report the Average Inference Time per Sequence (AITS) in seconds.
@@ 4.2 Comparison with Baseline Methods
Method FIDk($\downarrow$) FIDg($\downarrow$) Divk($\rightarrow$) Divg($\rightarrow$) MPJPE($\downarrow$) MPJVE($\downarrow$) BAS($\rightarrow$) FIDcd($\downarrow$) Divcd($\rightarrow$) BED($\rightarrow$) AITS($\downarrow$)
Ground Truth $0$- $0$- 10.86 7.82 0- - 0.1791 $00$- 12.53 0.5308 -
GestureLSM 14.65 034.23 12.20 09.00 171.37 19.01 0.1734 040.09 09.45 0.1903 3.13
EDGE 68.68 113.25 05.62 10.05 235.52 20.76 0.1854 1523.27 12.37 0.1904 2.91
TCDiff 105.97 251.35 14.23 24.54 182.64 22.91 0.2270 1472.00 11.31 0.2304 3.34
InterGen 35.89 047.28 12.24 09.13 210.66 18.69 0.1634 176.19 18.10 0.2746 5.22
Duolando 27.68 35.01 10.95 08.70 174.54 18.72 0.2086 017.49 14.73 0.3285 4.41
@@
Analysis. As reported in Table 1, our method, ReactDance, demonstrates superior performance across the majority of metrics. Note that these results are computed on the entire test set, which features long sequences with an average length of 2066 frames. This highlights our model’s strong generalization capability and consistency in long-term generation. The significant lead in realism ($\text{FID}_{k},\text{FID}_{g}$) and accuracy (MPJPE, MPJVE) stems from our effective HFSQ representation, which enables the diffusion model to learn motion hierarchically from coarse postural dynamics to fine-grained articulations. This high-fidelity representation directly enhances duet coherence, where ReactDance significantly outperforms all baselines (e.g., achieving a $\text{FID}_{cd}$ of 14.17 and a BED of 0.3863). Furthermore, ReactDance also achieves the most efficient generation with an AITS of 1.75s due to the parallel generation capability of our BLC sampling strategy. In contrast, competing methods exhibit key limitations. Architectures with a two-stage design like GestureLSM and Duolando, while competitive in solo motion quality, are constrained by their representations. GestureLSM’s use of a single-scale latent space for generation limits its ability to capture hierarchical interactions (evidenced by a high $\text{FID}_{cd}$ of 40.09), while Duolando’s VQ-VAE is less effective at fine-grained reconstruction. Other models like EDGE and TCDiff perform poorly overall, as they learn directly on the high-frequency raw motion space and lack explicit mechanisms for interaction modeling.
@@ 4.3 Ablation Study
Table 2: Ablation study of HFSQ and Progressive Masking (PM).
@@
w/o HFSQ $36.73$ $8.82$ $139.42$ $0.2040$
w/o PM $10.46$ $7.65$ 132.19 0.2003
@@
We ablate our Hierarchical Finite Scalar Quantizer (HFSQ) and Progressive Masking (PM) strategy in Table 2. First, replacing HFSQ with a comparable RVQ-VAE significantly degrades performance (FIDg increases from 7.63 to 36.73; MPJPE increases from 132.99 to 139.42), underscoring the superior motion representation capability of HFSQ. Second, removing the PM strategy trades a marginal MPJPE improvement for a substantial drop in realism (FIDg increases to 10.46), demonstrating that PM is crucial for synthesizing high-fidelity motion. Additional results are available in the supplementary materials.
@@
Table 3: Ablation of Dense Sliding Window (DSW) training and periodic sampling strategies (PPE & PCAM).
@@
We ablate the two core components of our Blockwise Local Context (BLC) sampling to validate its contributions to duet coherence and sampling efficiency, with results presented in Table 3. We analyze the dense sliding window component by increasing the training stride from 4 (our full model) to 16 and 64. We observe a significant degradation in coherence, as evidenced by the substantial increase in $\text{FID}_{cd}$ and the decline in BED. Moreover, we replace our periodic positional encoding (PPE) and causal attention masking (PCAM) with the latent stitching operation (Tseng et al., 2023). This change harms both sampling efficiency (AITS increases from 1.75 to 2.03) and generation quality ($\text{FID}_{cd}$ increases from 14.17 to 18.59). These results confirm a dense training stride is crucial for coherence, and our periodic sampling mechanism is key to achieving both high quality and efficiency.
Effect of Layer-Decoupled Guidance. Our diffusion model operates on hierarchical HFSQ latents, enabling us to apply independent classifier-free guidance weights $s_{r}$ of Eq. (9) to each motion scale. This provides fine-grained control over the fidelity-diversity trade-off across different semantic levels. For instance, a user can enforce strong postural alignment with high guidance on coarse scales while encouraging creative limb articulation with lower guidance on fine scales. A detailed analysis demonstrating this flexibility is available in the supplementary materials.
@@ 5 Conclusion
We present ReactDance, a diffusion framework for generating high-fidelity and coherent long-form reactive dance. Our method is built upon a hierarchical latent representation that captures multi-scale spatiotemporal dynamics. This representation further enables two key contributions: an efficient parallel sampling strategy for long sequence generation and a multi-level conditioning mechanism that offers fine-grained artistic control. ReactDance offers a comprehensive system for generating dynamic body motion under complex conditions, providing a versatile tool for music-driven character animation and digital choreography.
@@
Limitation. Despite the effectiveness of our HFSQ representation, it lacks explicit semantic meaning across scales, which reduces control interpretability. Additionally, finger motions are omitted mainly due to the noisy hand data of DD100 dataset, which are crucial for nuanced gestures. Future work will focus on imbuing our representation with semantic meaning for more intuitive control and incorporating detailed hand articulations to broaden the model’s expressive range.

### v3

@@ 4.1 DATASETS AND EXPERIMENTAL SETTING
Evaluation Metrics. We adopt a comprehensive set of quantitative metrics to evaluate the generated motion from four key perspectives: (1) Motion Quality. To assess the realism of individual motions, we measure distributional similarity and reconstruction accuracy. Distributional metrics include Fréchet Inception Distance (FID) and Diversity (Div) on both kinematic ($\text{FID}_{k},\text{Div}_{k}$(Onuma et al., 2008) and graphical ($\text{FID}_{g},\text{Div}_{g}$(Müller et al., 2005) features. Reconstruction accuracy against the ground truth is measured by Mean Per-Joint Position Error (MPJPE) for static poses and Mean Per-Joint Velocity Error (MPJVE) for dynamics (Ghosh et al., 2024). Additionally, the Physical Foot Contact (PFC) (Tseng et al., 2023) detects foot skating artifacts by verifying kinematic consistency. (2) Interaction Quality. To evaluate leader-reactor coordination, we compute FID and Diversity on cross-distance features ($\text{FID}_{cd},\text{Div}_{cd}$), which describe the evolving spatial relationships between the dancers’ key joints. To assess physical validity, we calculate the Inter-Penetration Rate (IPR), quantifying vertex-mesh penetrations using the Generalized Winding Number field (Jacobson et al., 2013), accelerated by AABB-based culling. (3) Music-Dance Alignment. We use two metrics: the Beat Align Score (BAS) (Siyao et al., 2022) to measure synchronization between motion peaks and music beats, and the Beat Echo Degree (BED) (Siyao et al., 2024) to quantify rhythmic consistency between the dancers. (4) Generation Efficiency. We report the Average Inference Time per Sequence (AITS) in seconds.
@@ 4.2 Comparison with Baseline Methods
Method FIDk($\downarrow$) FIDg($\downarrow$) Divk($\rightarrow$) Divg($\rightarrow$) MPJPE($\downarrow$) MPJVE($\downarrow$) PFC($\downarrow$) BAS($\rightarrow$) FIDcd($\downarrow$) Divcd($\rightarrow$) BED($\rightarrow$) IPR($\downarrow$) AITS($\downarrow$)
Ground Truth $0$- $0$- 10.86 7.82 0- - - 0.1791 $00$- 12.53 0.5308 - -
GestureLSM 14.65 034.23 12.20 09.00 171.37 19.01 0.7903 0.1734 040.09 09.45 0.1903 19.01 3.13
EDGE 68.68 113.25 05.62 10.05 235.52 20.76 0.6226 0.1854 1523.27 12.37 0.1904 08.44 2.91
TCDiff 105.97 251.35 14.23 24.54 182.64 22.91 1.1165 0.2270 1472.00 11.31 0.2304 07.58 3.34
InterGen 35.89 047.28 12.24 09.13 210.66 18.69 0.9842 0.1634 176.19 18.10 0.2746 17.58 5.22
Duolando 27.68 035.01 10.95 08.70 174.54 18.72 0.9276 0.2086 017.49 14.73 0.3285 17.42 4.41
@@
Analysis. As reported in Table 1, our method, ReactDance, demonstrates superior performance across the majority of metrics. Note that these results are computed on the entire test set, which features long sequences with an average length of 2066 frames. This highlights our model’s strong generalization capability and consistency in long-term generation. The significant lead in realism ($\text{FID}_{k},\text{FID}_{g}$) and accuracy (MPJPE, MPJVE) stems from our effective HFSQ representation, which enables the diffusion model to learn motion hierarchically from coarse postural dynamics to fine-grained articulations. This fine-grained capability also ensures high physical plausibility, as evidenced by the lowest PFC score (0.6039), indicating minimized foot skating artifacts compared to baselines. Regarding duet coherence, ReactDance significantly outperforms all baselines, achieving a $\text{FID}_{cd}$ of 14.17 and a BED of 0.3863. Crucially, our method maintains a low IPR (7.84%), significantly surpassing the autoregressive baseline Duolando (17.42%) and proving that our model generates spatially coherent interactions without severe collision artifacts. Furthermore, ReactDance also achieves the most efficient generation with an AITS of 1.75s due to the parallel generation capability of our BLC sampling strategy. In contrast, competing methods exhibit key limitations. Architectures with a two-stage design like GestureLSM and Duolando, while competitive in solo motion quality, are constrained by their representations. GestureLSM’s use of a single-scale latent space for generation limits its ability to capture hierarchical interactions (evidenced by a high $\text{FID}_{cd}$ of 40.09), while Duolando’s VQ-VAE is less effective at fine-grained reconstruction. Other models like EDGE and TCDiff perform poorly overall, as they learn directly on the high-frequency raw motion space and lack explicit mechanisms for interaction modeling.
@@ 4.3 Ablation Study
We conduct a comprehensive ablation study on the tokenizer architecture and Progressive Masking (PM) training strategy in Table 2. Comparing different architectures, the VAE suffer from posterior collapse, leading to over-smoothed reconstruction (MPJPE 40.87) and poor generation alignment
Table 2: Ablation study of HFSQ tokenizer and the Progressive Masking (PM).
@@
Stage 1: Reconstruction
VAE $5.32$ 7.78 $40.87$ 0.1805
RVQ-VAE (w/o. PM) $4.57$ $7.72$ $36.59$ $0.1841$
RVQ-VAE (w PM) $4.09$ $7.73$ $32.98$ $0.1842$
HFSQ (w/o. PM) 3.77 7.74 24.15 0.1784
HFSQ (w PM, Ours) 3.56 $7.73$ 28.66 $0.1816$
Stage 2: Generation
VAE $18.99$ $7.77$ $230.21$ $0.2705$
RVQ-VAE (w/o. PM) $36.73$ $8.82$ $139.42$ $0.2040$
RVQ-VAE (w PM) $26.98$ 7.75 $138.28$ 0.1946
HFSQ (w/o. PM) 10.46 $7.65$ 132.19 0.2003
@@
(MPJPE 230.21). Standard RVQ-VAE with same number of residual layers ($R=2$) improves upon VAE but still lags behind HFSQ, particularly in generation quality (FIDg 26.98 vs. 7.63), even when trained with PM strategy. We attribute this to the irregular topology of vector codebooks, where the latent of adjacent indices lack semantic continuity, creating a difficult optimization landscape for diffusion. In contrast, HFSQ projects motion onto a fixed scalar grid that preserves ordinal relations, providing a smoother, “diffusion-friendly” manifold. Finally, analyzing the PM strategy reveals a critical trade-off: removing the PM strategy trades a marginal MPJPE improvement for a substantial drop in realism (FIDg increases to 10.46), demonstrating that PM is crucial for synthesizing high-fidelity motion. Additional results are available in the Appendix A.3.
@@
Table 3: Ablation of Dense Sliding Window (DSW) training and BLC sampling strategies.
@@
We ablate the two core components of our Blockwise Local Context (BLC) sampling to validate its contributions to duet coherence and sampling efficiency, with results presented in Table 3. We analyze the dense sliding window component by increasing the training stride from 4 (our full model) to 16 and 64. We observe a significant degradation in coherence, as evidenced by the substantial increase in $\text{FID}_{cd}$ and the decline in BED. Moreover, we replace our periodic positional encoding (PPE) and periodic causal attention masking (PCAM) of BLC with the latent stitching operation (Tseng et al., 2023). This change harms both sampling efficiency (AITS increases from 1.75 to 2.03) and generation quality ($\text{FID}_{cd}$ increases from 14.17 to 18.59). These results confirm a dense training stride is crucial for coherence, and our periodic sampling mechanism is key to achieving both high quality and efficiency. For a comprehensive analysis of the trade-off between motion continuity and training cost, please refer to Appendix A.7.
Effect of Layer-Decoupled Guidance. Our diffusion model operates on hierarchical HFSQ latents, enabling us to apply independent classifier-free guidance weights $s_{r}$ of Eq. (9) to each motion scale. This provides fine-grained control over the fidelity-diversity trade-off across different semantic levels. For instance, a user can enforce strong postural alignment with high guidance on coarse scales while encouraging creative local articulation with lower guidance on fine scales. A detailed analysis demonstrating this flexibility is available in Appendix A.4.
@@ 5 Conclusion
We present ReactDance, a diffusion framework for generating high-fidelity and coherent long-form reactive dance. Our method is built upon a hierarchical latent representation that captures multi-scale spatiotemporal dynamics. This representation enables two key contributions: an efficient parallel sampling strategy for long sequence generation and a multi-level conditioning mechanism that offers fine-grained artistic control. By solving the fundamental challenge of maintaining kinematic stability over extended durations, ReactDance serves as a robust cornerstone for coherent long-form reactive dance generation. We hope this foundation paves the way for future research to ascend from physical continuity to high-level semantic modeling, enabling the exploration of narrative-driven choreography and emotional storytelling in reactive dance.
@@
Limitation. Despite the effectiveness of our HFSQ representation, it lacks explicit semantic meaning across scales, which reduces control interpretability. Additionally, finger motions are omitted mainly due to the noisy hand data of DD100 dataset, which are crucial for nuanced gestures.
Acknowledgements. We sincerely acknowledge the anonymous reviewers for their insightful feedback. This work is partially supported by the Natural Science Foundation of China (Grant No. 62472178), the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems (No. MAIS2024111), the Fundamental Research Funds for the Central Universities, and the Science and Technology Commission of Shanghai Municipality (Grant No. 25511107200).
Appendix A Appendix
In the appendix, we present:
  • Section A.1: The Use of Large Language Models (LLMs).
  • Section A.2: Training Hyper-parameters.
  • Section A.3: More Results of Ablation on HFSQ Representation and Progressive Masking.
  • Section A.4: More Results of Layer-Decoupled Guidance.
  • Section A.5: User Study Results.
  • Section A.6: More Results on Residual Stages $R$ and Scalability Analysis.
  • Section A.7: Additional Analysis on DSW Stride Selection.
  • Section A.8: Hand Motion Modeling Analysis.
  • Section A.9: Ablation of Conditioning Signals.
  • Section A.10: Generalization via Role Swapping.
  • Section A.11: Zero-Shot Generalization on Out-of-Distribution Data.
A.1 The USE OF LARGE LANGUAGE MODELS
In the preparation of this manuscript and its accompanying code, we utilized Large Language Models (LLMs) as assistive tools. We wish to clarify that the core scientific contributions—including the problem formulation, the design of our proposed method, and the experimental analysis—are entirely the work of the authors. The use of LLMs was confined to the following supporting tasks:
  • Manuscript Preparation: LLMs were employed for language refinement. This included improving grammatical correctness, rephrasing sentences for better clarity and flow, and ensuring stylistic consistency throughout the paper. All final text was carefully reviewed and edited by the authors to ensure it accurately reflects our research and contributions.
  • Software Development: LLMs assisted in writing auxiliary and boilerplate code. Specific use cases included generating scripts for data preprocessing and visualization, creating utility functions, and debugging isolated code segments. The core algorithmic implementation of our proposed method was developed exclusively by the authors.
In all instances, the LLM served as a productivity and quality-enhancement tool. The authors maintain full intellectual responsibility and ownership for all conceptual and experimental aspects of this work.
A.2 TRAINING HYPER-PARAMETERS
Our framework is implemented in PyTorch and trained on a single NVIDIA RTX L40 GPU. The motion HFSQ is trained for 200 epochs using the Lion optimizer (Chen et al., 2023) with initial learning rate $3e^{-5}$ and a stepwise learning rate scheduler decays the rate by 0.998 per epoch. For diffusion training, the denoise network is trained for 1500 epochs using the Adamw optimizer with initial learning rate $2e^{-4}$ and a stepwise learning rate scheduler decays the rate by 0.998 per epoch. Batch size is set to 256 across all stages.
A.3 MORE RESULTS OF ABLATION ON HFSQ REPRESENTATION AND PROGRESSIVE MASKING
The visualization in Fig. 6 supports the quantitative findings in the main paper. The ablated model using an RVQ-VAE often generates poses that are individually plausible but contextually incorrect within the duet. This leads to clear visual artifacts such as inaccurate relative positioning and distance between the dancers, as well as failures in fine-grained hand-to-hand or hand-to-body contact. This demonstrates that HFSQ’s hierarchical structure is crucial for effectively representing the complex spatial dynamics of dance interactions.
Figure 6: Qualitative comparison for the HFSQ ablation. While our full model maintains correct spatial relationships, the model trained without HFSQ (using an RVQ-VAE) fails to capture precise inter-personal dynamics, resulting in incorrect relative distancing and mismatched hand interactions (highlighted in red circles).
As shown in Fig. 7, removing the PM strategy introduces severe interaction artifacts. This aligns with the quantitative trade-off noted in the main paper, where the ablated model achieves a slightly lower MPJPE but a much worse FIDg. Without the regularization provided by PM, the model may reconstruct individual poses more closely but fails to learn robust interaction context. This leads to fundamental errors in character orientation relative to the leader and results in physically implausible outcomes, such as body part interpenetration and collisions. This highlights that the PM strategy is essential for learning the physical constraints and relational logic of reactive dance.
Figure 7: Qualitative comparison for the Progressive Masking (PM) ablation. Our full model generates a coherent and physically plausible interaction. In contrast, removing the PM strategy leads to critical errors in interaction logic, such as the reactor adopting an incorrect orientation relative to the leader, resulting in interpenetration artifacts (highlighted).
A.4 MORE DETAILS ABOUT LAYER-DECOUPLED GUIDANCE
Our Layer-Decoupled Classifier-Free Guidance (LDCFG) is designed to overcome the limitation of a single, global guidance weight used in conventional CFG. By assigning independent guidance strengths $s_{r}$ (Eq. 9) to each hierarchical layer of our HFSQ representation, LDCFG provides fine-grained control over the fidelity-diversity trade-off at different semantic levels.
We conducted a quantitative analysis by sampling with different LDCFG weight pairs $S=[s_{\text{coarse}},s_{\text{fine}}]$, where $s_{\text{coarse}}$ and $s_{\text{fine}}$ control the coarse- and fine-grained motion scales, respectively. The results are presented in Table 4. The data reveals a clear trade-off between solo motion fidelity and interaction quality. For instance, increasing the guidance weights (e.g., $S=[1.2,1.5]$) improves reconstruction metrics like MPJPE and MPJVE, but at the cost of harming the interaction quality (FIDcd increases significantly to 18.57). Conversely, high guidance on both scales ($S=[1.5,1.5]$) improves diversity (Divk, Divg) but leads to the worst interaction scores (FIDcd of 24.56). Our chosen setting, $S=[1.2,1.2]$ (highlighted), provides an optimal balance, achieving the best FIDk and strong reconstruction quality while maintaining competitive interactive metrics. This demonstrates that LDCFG enables a nuanced tuning of realism against interaction coherence.
Furthermore, because our representation disentangles body components, we can apply LDCFG for targeted artistic control. Fig. 8 visualizes this capability, showcasing independent guidance applied to the upper body, lower body, and their combination to create distinct stylistic motion variations while maintaining plausible interactions.
Table 4: Solo and Duet metrics with diverse Layer-Decoupled Classifier-Free Guidance (LDCFG) weights. Best and second-best results are in bold and underline, respectively.
Solo Metrics Interactive Metrics
LDCFG Weight FIDk($\downarrow$) FIDg($\downarrow$) Divk($\rightarrow$) Divg($\rightarrow$) MPJPE($\downarrow$) MPJVE($\downarrow$) BAS($\rightarrow$) FIDcd($\downarrow$) Divcd($\rightarrow$) BED($\rightarrow$)
Ground Truth $10.86$ $7.82$ $0.1791$ $12.53$ $0.5308$
$S=[1.0,1.0]$ $8.09$ $10.37$ $10.74$ $7.69$ $155.30$ $17.59$ $0.2198$ $18.15$ $9.82$ $0.3470$
$S=[1.0,1.2]$ $6.17$ $8.10$ $10.79$ $7.74$ $138.68$ $16.29$ $0.2154$ 13.66 $10.55$ $0.3709$
$S=[1.0,1.5]$ $5.77$ $7.70$ $10.80$ $7.75$ $134.78$ $15.90$ $0.2091$ 13.40 10.65 $0.3835$
[…]
$S=[1.2,1.5]$ 5.62 7.29 $10.83$ $7.79$ 132.20 15.43 0.1978 $18.57$ $10.14$ 0.3901
$S=[1.5,1.5]$ $6.07$ 7.62 10.85 7.81 $135.27$ 15.62 0.1961 $24.56$ $9.68$ 0.3925
Figure 8: Qualitative visualization of LDCFG’s body-part control. We apply decoupled guidance with varying strengths to different body regions. The blue robot is the leader. The other colored robots are generated by our model, each with a different LDCFG setting applied to the (a) lower body, (b) upper body, (c) their combination, or (d) full body global motion demonstrating fine-grained artistic control.
A.5 User Study Results
Figure 9: User study results. ReactDance consistently outperforms all baselines across all criterions.
A.6 More Results on Residual Stages $R$ and Scalability Analysis
In this section, we provide a detailed scalability analysis regarding the depth of our Hierarchical Finite Scalar Quantization (HFSQ), specifically the number of residual stages $R$. The choice of $R$ involves a critical trade-off between the representational capacity of the tokenizer and the modeling complexity for the subsequent diffusion generator.
We evaluate the performance across $R\in\{1,2,3,4\}$ in Table 5. The results reveal three key observations:
  • Reconstruction Fidelity: Increasing $R$ from 1 to 2 yields a substantial reduction in reconstruction error (MPJPE decreases from $35.53$ to $24.32$), indicating that the second stage is essential for capturing fine-grained articulation. However, further increasing $R$ to 3 or 4 provides diminishing returns (MPJPE only improves marginally to $22.97$) while linearly increasing the latent space size.
  • Generation Quality: Surprisingly, higher stages ($R\geq 3$) lead to a degradation in generation quality (FIDk increases from $5.57$ to $6.10$). We attribute this to the increased complexity of the latent distribution; as the sequence of residual codes becomes longer, the joint distribution becomes harder for the diffusion model to learn effectively given the same training budget.
  • Computational Cost: The training time increases only marginally with $R$, rather than scaling proportionally. Specifically, $R=2$ incurs a negligible overhead compared to $R=1$ (1.00x vs. 0.94x) while offering significantly better motion quality.
Based on this analysis, $R=2$ represents the optimal performance-efficiency trade-off, achieving the best balance between reconstruction fidelity, generative realism, and computational efficiency. Consequently, we adopt $R=2$ as our default setting.
Table 5: Scalability Analysis of Residual Stages $R$. While higher $R$ marginally improves reconstruction, $R=2$ achieves the best generation quality (FID) and the optimal trade-off between performance and cost.
Reconstruction Metrics (HFSQ) Generation Metrics (Diffusion) Training Cost
[…]
$R=1$ 2.69 35.53 13.67 13.10 10.15 24.50 0.94x
[…]
$R=3$ 0.38 24.08 2.35 5.92 7.85 14.50 1.07x
$R=4$ 0.30 22.97 2.10 6.10 8.02 14.85 1.12x
A.7 Additional Analysis on DSW Stride Selection
In Section 4.3, we validated the impact of training stride $s$ on interaction coherence. Here, we extend this ablation to include Motion Continuity (measured by Boundary Jitter) and Training Cost (relative training time), providing a justification for our design choice.
We report the comparison across a wide range of strides in Table 6. We define Jitter as the mean L2 velocity change at block boundaries (half-radius of 30 frames) to quantify the smoothness of the stitched motions.
Analysis of Motion Continuity. As the stride $s$ decreases (density increases), the model receives more frequent supervision on boundary transitions. Consequently, the Jitter score consistently improves ($18.08\to 15.19$), confirming that the Dense Sliding Window (DSW) mechanism effectively acts as a phase-agnostic kinematic smoother. Notably, sparse strides ($s\geq 64$) fail to capture these transitions, resulting in severe discontinuities (>17.9>17.9) and collapsed interaction quality ($\text{FID}_{cd}>39$).
[…]
Table 6: Full Ablation of Dense Sliding Window Stride ($s$). Jitter measures boundary discontinuity.
Motion Continuity Interaction Quality Training Cost
DSW Stride Setting Jitter ($\downarrow$) FIDcd ($\downarrow$) Train Time ($\downarrow$)
[…]
A.8 Hand Motion Modeling Analysis
The DD100 dataset contains inherent high-frequency noise and jitter in the ground truth hand motions (Noisy DD100 GT_Hand Video), and conventional models often overfit to this noise during training, resulting in unstable generation. While baseline methods struggle to generate coherent reactive dances even without hand motions (see Duolando Failure Video), the inclusion of fine-grained hand modeling further exacerbates their performance degradation.
In contrast, our investigation reveals that ReactDance acts as a robust motion denoiser. This capability stems from the quantization mechanism of our Hierarchical Finite Scalar Quantization (HFSQ). By projecting continuous motion into discrete scalar codebooks, HFSQ effectively filters out high-frequency artifacts while preserving meaningful articulation. As shown in Table 7, ReactDance achieves significantly lower Jitter (25.32) compared to the autoregressive baseline Duolando (32.04). Visual comparisons are provided in the Hand Modeling Comparison Video.
Table 7: Hand Motion Analysis. ReactDance effectively filters out high-frequency artifacts inherent in the dataset, significantly outperforming Duolando in motion stability (lower Jitter).
Method MPJPE $\downarrow$ Jitter $\downarrow$ FIDg $\downarrow$
Duolando 340.40 32.04 9.37
[…]
A.9 Ablation of Conditioning Signals
We investigate the distinct roles of the leader’s motion and the musical accompaniment in guiding the generation process. We trained variants of our model by removing each condition individually. The results are summarized in Table 8.
  • Effect of Leader Motion: Removing the leader’s motion causes a collapse in interaction quality, as evidenced by the sharp degradation in FIDcd ($14.2\to 48.8$). This confirms that the leader’s signal is critical for establishing the structural and spatial framework of the reaction.
  • Effect of Music: Interestingly, removing the music condition leads to a slight improvement in MPJPE (reconstruction error) but a significant degradation in interaction realism (FIDcd $14.2\to 35.4$). This phenomenon occurs because, without music, the model becomes a “mechanical tracker”, tending to generate safe poses to minimize geometric error. However, it loses the nuances required for high-fidelity dance generation, confirming that music is essential for fine-grained interactions.
Table 8: Ablation of Conditioning Inputs. Results indicate that the leader’s motion is crucial for structural interaction, while music is essential for fine-grained interactive realism.
Solo Metrics Interactive Metrics
Method MPJPE $\downarrow$ MPJVE $\downarrow$ FIDcd $\downarrow$ BAS $\rightarrow$
Ground Truth - - - 0.1791
w/o. Leader Motion 238.50 22.76 48.82 0.2074
w/o. Music 130.27 15.05 35.35 0.1947
[…]
A.10 Generalization via Role Swapping
To verify that ReactDance learns generalized interaction rules rather than memorizing fixed leader-follower pairs, we conducted a Role Swapping experiment. Specifically, we inverted input conditions by feeding follower’s motion as the leader signal and tasked the model to generate the leader’s part.
As demonstrated in Role Swapping Video, the model successfully generates valid, semantically coherent reactions for the reversed pairs. This confirms that ReactDance captures the underlying physical and choreographic rules of interaction (e.g., synchronization, relative spacing) independent of specific character roles.
A.11 Zero-Shot Generalization on Out-of-Distribution Data
To assess the robustness of ReactDance beyond the specific choreography of the training set (DD100), we conducted a Zero-Shot evaluation using the FineDance dataset (Li et al., 2023). FineDance contains diverse solo dance genres that are disjoint from our training data, serving as rigorous Out-of-Distribution (OOD) input for the leader.
Despite the significant domain shift in dance styles and musical accompaniment, ReactDance synthesizes physically plausible and semantically coherent reactions. This suggests that the model has successfully learned generalized interaction dynamics—such as velocity matching, spatial awareness, and rhythmic synchronization—rather than merely memorizing dataset-specific motion patterns.
Qualitative demonstrations of this generalization capability are provided in following videos: